# A theory of phonon-induced friction on molecular adsorbates

### v2

@@ Abstract
In this manuscript, we provide a general theory for how surface phonons couple to molecular adsorbates. Our theory maps the extended dynamics of a solid’s atomic vibrational motions to a generalized Langevin equation, and by doing so captures these dynamics in a single quantity: the non-Markovian friction. The different frequency components of this friction are the phonon modes of the solid weighted by their coupling to the adsorbate degrees of freedom. Using this formalism, we demonstrate that physisorbed species couple primarily to acoustic phonons while chemisorbed species couple to dispersionless local vibrations. We subsequently derive equations for phonon-adjusted reaction rates using transition state theory and demonstrate that these corrections improve agreement with experimental results for CO desorption rates from Pt(111).
@@ 1 Introduction
Our theoretical model describes the motion of surface-bound molecules via a generalized Langevin equation (GLE). The central parameter of this equation is the non-Markovian friction kernel (or memory kernel), which encapsulates the different surface vibrational modes weighted by their coupling to the adsorbate. Using this formalism, we demonstrate that the influence of surface phonons on the dynamics of an adsorbed molecule depends significantly on the magnitude of the adsorbate-surface coupling. Specifically, we find that chemisorbed species couple primarily to dispersionless local vibrations, while physisorbed species couple primarily to acoustic phonons across a broad range of frequencies. The key parameter that determines whether the primary influence of surface vibrations is via extended or localized phonon modes is the ratio of the solid’s Debye frequency to the frequency of the adsorbate-surface bond (Figure 1). By combining these observations with harmonic transition state theory, we derive equations that describe how phonons alter reaction rates at solid surfaces and demonstrate that these phononic corrections agree with experimental measurements of desorption rate constants.
@@
The approach we present in this manuscript extends our previous method for utilizing the GLE for describing the influence of surface vibrational modes on molecular adsorbates. farahvash_using_2023 In our previous method, the GLE was designed to describe the collective influences of the surrounding solid on a single representative surface atom (i.e., binding site). In the method we present here, the GLE is designed to describe the influence of the entire surface on the dynamics of the adsorbate itself. This new approach is more easily interpretable because the friction kernel directly represents influence of surface phonons on the adsorbate.
@@ 2.1 Theory
In this section, we derive a theory for modeling the effects of solid surface vibrations on the dynamics of an adsorbate. The general approach involves integrating out the nuclear degrees of freedom of the solid and representing their influence as effective equations of motion for the adsorbate. In order to do so we must make two critical assumptions. First, we assume that the nuclear degrees of freedom of the adsorbate and solid can be described classically and adiabatically. There are cases where adsorbate dynamics are non-adiabatic, and in such cases a mixed quantum-classical GLE approach has been shown to be effective 27, 28, however the development and analysis of such a theory is beyond the scope of this paper. Second, we assume that the solid potential energy surface and surface-adsorbate interaction are both harmonic. This approximation simplifies the formulation and is appropriate for the development of generalizable physical insight. We note that the effects of anharmonicity can be incorporated into the theory, which would be a natural target of future developments.
@@
where $\mathbf{C}=\mathbf{G}_{\mathrm{AS}}\mathbf{U}$ is a matrix of couplings between each adsorbate coordinate and solid phonon mode, and $\boldsymbol{\omega}^{2}=\mathbf{U}^{T}\mathbf{H}_{\mathrm{S}}\mathbf{U}$ is a diagonal matrix of the squared frequencies of the solid phonons. Combining Eq. 8 and Eq. 9 yields a GLE for the adsorbate degrees of freedom29, 30,
@@
In some cases Eqs. 11 and 13 may further be simplified by stating that adsorbates couple only to a handful of surface sites and thus the matrices $\mathbf{G}_{AB}$ and $\mathbf{C}$ are sparse. For example, consider the case of an adsorbate that interacts with a single surface atom via a single adsorbate atom (e.g. CO on Pt(111) 31, 32). If the force constant associated with this interaction is $\mu\omega_{\mathrm{as}}^{2}$, where $\mu$ is the reduced mass of the adsorbate and surface atom, Eqs. 11 and 13 may be simplified to,
@@ 2.2 Results and discussion
We have computed the phonon friction kernel for a simple model of CO on Pt(111). The CO was modeled as a single adatom and assumed the interact with a single adsorption site, as experimental structures show that CO adsorbs primarily atop Pt(111) sites 31, 32. We artificially vary the adsorbate-surface frequency $\omega_{\mathrm{as}}$ over a range of physically motivated values, in order asses how the strength of the adsorbate-surface interaction affects the properties of the friction kernel. The metal potential energy surface $V_{S}$ was modeled using an effective medium theory (EMT) forcefield developed by Norskov et al.33 More details about calculations can be found in the Section 5.
@@
Figure 2 illustrates the results of our calculations for a selected set of $\omega_{\mathrm{as}}$ values ranging from those characteristic of weakly adsorbed species to those characteristic of strongly adsorbed species. We plot the friction kernel computed in the direction normal to the surface in systems with either a 4x4x8 or 8x8x8 Pt(111) slab. The lower values of $\omega_{\mathrm{as}}$ that we consider lie below that of the platinum’s Debye frequency of $\omega_{\mathrm{D}}=156\mathrm{cm}^{-1}$. We associate these values with the physisorbed regime, as they are characteristic of weakly adsorbed systems such as noble gases on Pt(111).34 The higher values of $\omega_{\mathrm{as}}$ that we consider are associated with the chemisorbed regime, corresponding to physical systems such as CO on Pt(111), which has $\omega_{\mathrm{as}}\approx 480\ \mathrm{cm}^{-1}$).32, 35, 36 Delta function peaks in $\bar{K}(\omega)$ and $\rho(\omega)$ were broadened to thin Lorentzians of width $1\mathrm{cm}^{-1}$ for ease of visualization.
@@
For all cases of $\omega_{\mathrm{as}}$, the adsorbate couples strongly to a surface acoustic mode appearing around $\omega=10\mathrm{cm}^{-1}$. The location of this peak is highly dependent on the dimensions of the simulated surface slab, consistent with the behavior of an acoustic phonon. In Sec. 3, we give a more detailed theoretical and numerical analysis of the size effects observed here, fully accounting for phonon dispersion. In the physisorbed regime, the friction kernel is dominated by this acoustic mode, resulting in highly non-Markovian behavior. The phonon mode associated with this peak is associated with the flexing of the lattice in the direction perpendicular to the surface. Snapshots of this motion are provided in the insets in Figure 1 and video animations are provided in the electronic SI.
@@
In the chemisorbed regime, the friction kernel is dominated by a high frequency mode around $\omega=210\ \mathrm{cm}^{-1}$. This mode arises from the local oscillations of the surface site the adsorbate is bound to. The frequency of this mode is independent of simulation cell size, suggesting that it is dispersionless. The large amplitude of this mode signifies that in the chemisorbed limit, the adsorbate is primarily sensitive to the local oscillations of the surface binding site (which are shifted in frequency due to the presence of the adsorbate). Many published models for reactive scattering on solid surfaces will describe solid vibrations using only a single, harmonically bound surface atom 6, 37, 38, 39, 40. These results explain why such a method is successful for strongly coupled species.
@@
Despite this large shift in the spectral density when varying $\omega_{\mathrm{as}}$, in Figure 2C we demonstrate that the phonon density of states is nearly identical to that of a bare surface. The only significant change in $\rho(\omega)$ is the presence of the aforementioned surface site local mode present around $\omega=210\ \mathrm{cm}^{-1}$. These results make physical sense given that the low frequency acoustic modes of a solid should be unaffected by gaseous species, especially at low pressures and surface coverages.
@@
To evaluate the effect of the solid force field and structure on these results, we have tested several different crystal structures, facets, elemental compositions, and force fields. The results of these tests are presented in the Section S1 of the SI. While the quantitative properties of the friction kernel vary across different systems (e.g. Ru has a much higher Debye frequency than Pt), the qualitative dependence of the friction kernel on $\omega_{\mathrm{as}}$ is quite general. This independence with respect to the details of the atomistic model is appealing and can be understood from a perturbative perspective. Specifically, in the physisorbed limit, the motion of the surface binding site is a small perturbation to the bulk phonon modes, while in the chemisorbed limit, the bulk phonon modes are a small perturbation to the motion of the surface binding sites. In the SI, we rigorously examine this statement by comparing perturbative schemes to the exact results for $K(t)$ and $\bar{K}(\omega)$.
@@
Finally, we note that in this section we have only studied the effects of phonons for an ideal clean surface. Presumably, sources of surface heterogeneity, such as steps and defects, may add further richness and complexity to the picture we have provided here. We also note that we have only studied elemental solids, and the optical modes of polyatomic crystals could add another interesting dimension to this physical picture.
@@ 3 Phonon dispersion
In a surface slab, the presence of anisotropy breaks the symmetry of the 3D crystal. This symmetry breaking leads to additional surface mode bands, which can be acoustic (Rayleigh waves), or non-acoustic in character41, 42, 43. In Fig. 4B. and 4C. we demonstrate the phonon dispersion of a Pt(111) surface calculated via EMT. These results were calculated using a 4x4x8 surface replicated in a 6x6 super-cell. We verified convergence with respect to supercell size. Of course, the number of surface phonon bands depends on the size of the surface unit cell used in computing $\mathbf{D}(\mathbf{k})$, however, we demonstrate in in Figure S7 results are qualitatively similar across different unit cell sizes.
@@
Figure 4: Friction kernel for a model adsorbate on Pt(111) in a finite system (black) and in the infinite system size limit (blue). The finite system is equivalent to the 4x4x8 results presented in Figure 1.
@@ 4 Phonon effects on reaction rates
One of the primary motivations for studying the atomic vibrations of a solid surface is to elucidate the role of phonons in reactions rates at catalytic interfaces. As was noted first by Kramers44, and expanded on by many others45, 46, 47, 48, the friction of the environment plays a critical role in determining reaction rates, and especially in determining rate prefactors. In this section, we use transition state theory (TST) to demonstrate how surface phonons affect reaction rates, emphasizing the role of the phonon friction kernel. We then specialize to the case of desorption rates and compare the results of our model to experimental temperature dependent rate constants for CO and Xe desorption from Pt(111). Several details of the formalism given this section have been omitted for the purpose of providing a concise presentation of the theory that focuses on physical ramifications. A thorough presentation of the theory is provided in the SI for interested readers.
@@ 4.1.1 General theory
where $N$ is the total number of modes, $f_{i}$ are the vibrational frequencies in the reactant basin, $f_{i}^{{\ddagger}}$ are the non-imaginary vibrational frequencies at the transition state. Note that since the transition state is defined to be a saddle point, it always has one imaginary frequency mode and this mode corresponds the reaction coordinate. Pollak illustrated that Eq. 22 is equivalent to Kramers and Grote-Hynes rate theories49. In a gas phase molecular reaction, $f_{i}$ is the $i$th eigenfrequency of the molecular mass-weighted Hessian. However, in a reaction at a surface interface, both surface phonon modes and molecular/adsorbate modes will contribute to the rate constant. We thus separate the product of eigenfrequencies into contributions from the solid surface and the molecular adsorbate to yield,
@@
Comparing Eqns. 22 and 23 reveals that phonons introduce two distinct corrections to rate constant: the ratio of solid phonon frequencies between the reactant and transition state, and the shift in adsorbate frequencies. The ratio of phonon frequencies is largest when the reactant molecules are strongly coupled to the surface and the transition state is not. In Figure 2C we demonstrated that an adsorbate strongly coupled to a surface affects only the highest frequency phonon mode, leaving the bulk of the phonon density of states unchanged. Thus, we can approximate the ratio of phonon frequencies as,
@@
where $\tilde{\omega}_{\mathrm{D}}$ is the highest frequency phonon mode when the reactants are bound to the surface and $\omega_{\mathrm{D}}$ is the bare solid Debye frequency.
@@
$\Delta\tilde{S}^{\ddagger}=k_{\mathrm{B}}\sum^{N_{A}-1}_{i=0}\ln\left(\frac{\tilde{f}^{{\ddagger}}_{i}}{\tilde{f}_{i}}\right).$ (27)

### v3

@@ Abstract
In this manuscript, we provide a general theory for how surface phonons couple to molecular adsorbates. Our theory maps the extended dynamics of a surface’s atomic vibrational motions to a generalized Langevin equation, and by doing so captures these dynamics in a single quantity: the non-Markovian friction. The different frequency components of this friction are the phonon modes of the surface slab weighted by their coupling to the adsorbate degrees of freedom. Using this formalism, we demonstrate that physisorbed species couple primarily to acoustic phonons while chemisorbed species couple to dispersionless local vibrations. We subsequently derive equations for phonon-adjusted reaction rates using transition state theory and demonstrate that these corrections improve agreement with experimental results for CO desorption rates from Pt(111).
@@ 1 Introduction
Our theoretical model describes the motion of surface-bound molecules via a generalized Langevin equation (GLE). The central parameter of this equation is the non-Markovian friction kernel (or memory kernel), which encapsulates the different surface vibrational modes weighted by their coupling to the adsorbate. This friction kernel can be developed to represent surface modes with wavelengths well beyond that of a typical simulation cell, thereby eliminating a common source of finite-size effects. Using this formalism, we demonstrate that the influence of surface phonons on the dynamics of an adsorbed molecule depends significantly on the magnitude of the adsorbate-surface coupling. Specifically, we find that chemisorbed species couple primarily to dispersionless local vibrations, while physisorbed species couple primarily to acoustic phonons across a broad range of frequencies. The key parameter that determines whether the primary influence of surface vibrations is via extended or localized phonon modes is the ratio of the solid’s Debye frequency to the frequency of the adsorbate-surface bond (Fig. 1). By combining these observations with harmonic transition state theory, we derive equations that describe how phonons alter reaction rates at solid surfaces, and demonstrate that these phononic corrections agree with experimental measurements of desorption rate constants.
The approach we present in this manuscript extends a previous method for utilizing the GLE for describing the influence of surface vibrational modes on molecular adsorbates. 27, 28 In the previous method, the GLE was designed to describe the collective influences of the surrounding solid on a single representative surface atom (i.e., binding site). In the method we present here, the GLE is designed to describe the influence of the entire surface on the dynamics of the adsorbate itself. This new approach is more easily interpretable because the friction kernel directly represents influence of surface phonons on the adsorbate.
@@ 2.1 Theory
In this section, we derive a theory for modeling the effects of surface vibrations on the dynamics of an adsorbate. The general approach involves integrating out the nuclear degrees of freedom of the solid and representing their influence as effective equations of motion for the adsorbate. In order to do so we must make two critical assumptions. First, we assume that the nuclear degrees of freedom of the adsorbate and solid can be described classically and adiabatically. There are cases where adsorbate dynamics are non-adiabatic, and in such cases a mixed quantum-classical GLE approach has been shown to be effective 29, 30, however the development and analysis of such a theory is beyond the scope of this paper. Second, we assume that the solid potential energy surface and surface-adsorbate interaction are both harmonic. This approximation simplifies the formulation of the theory and is appropriate for the development of generalizable physical insight. We note that the effects of anharmonicity can be incorporated into the theory, which would be a natural target of future developments.
@@
where $\mathbf{C}=\mathbf{G}_{\mathrm{AS}}\mathbf{U}$ is a matrix of couplings between each adsorbate coordinate and solid phonon mode, and $\boldsymbol{\omega}^{2}=\mathbf{U}^{T}\mathbf{H}_{\mathrm{S}}\mathbf{U}$ is a diagonal matrix of the squared frequencies of the solid phonons. Combining Eq. 8 and Eq. 9 yields a GLE for the adsorbate degrees of freedom31, 32,
@@
In some cases Eqs. 11 and 13 may further be simplified by stating that adsorbates couple only to a handful of surface sites and thus the matrices $\mathbf{G}_{AB}$ and $\mathbf{C}$ are sparse. For example, consider the case of an adsorbate that interacts with a single surface atom via a single adsorbate atom (e.g. CO on Pt(111) 33, 34). If the force constant associated with this interaction is $\mu\omega_{\mathrm{as}}^{2}$, where $\mu$ is the reduced mass of the adsorbate and surface atom, Eqs. 11 and 13 may be simplified to,
@@ 2.2 Results and discussion
We have computed the phonon friction kernel for a simple model of CO on Pt(111). The CO was modeled as a single adatom and assumed the interact with a single adsorption site, as experimental structures show that CO adsorbs primarily atop Pt(111) sites 33, 34. We artificially vary the adsorbate-surface frequency $\omega_{\mathrm{as}}$ over a range of physically motivated values, in order asses how the strength of the adsorbate-surface interaction affects the properties of the friction kernel. The metal potential energy surface $V_{S}$ was modeled using an effective medium theory (EMT) forcefield developed by Norskov et al.35 More details about calculations can be found in the Section 5.
@@
Fig. 2 illustrates the results of our calculations for a selected set of $\omega_{\mathrm{as}}$ values ranging from those characteristic of weakly adsorbed species to those characteristic of strongly adsorbed species. We plot the friction kernel computed in the direction normal to the surface in systems with either a 4x4x8 or 8x8x8 Pt(111) slab. The lower values of $\omega_{\mathrm{as}}$ that we consider lie below that of the platinum’s Debye frequency of $\omega_{\mathrm{D}}=156\mathrm{cm}^{-1}$. We associate these values with the physisorbed regime, as they are characteristic of weakly adsorbed systems such as noble gases on Pt(111).36 The higher values of $\omega_{\mathrm{as}}$ that we consider are associated with the chemisorbed regime, corresponding to physical systems such as CO on Pt(111), which has $\omega_{\mathrm{as}}\approx 480\ \mathrm{cm}^{-1}$).34, 37, 38 Delta function peaks in $\bar{K}(\omega)$ and $\rho(\omega)$ were broadened to thin Lorentzians of width $1\mathrm{cm}^{-1}$ for ease of visualization. It is important to note that the non-decaying oscillations in $K(t)$ are a consequence of the harmonic approximation. Accounting for anharmonicities will cause the memory kernel to asymptotically decay at long-times.
@@
For all cases of $\omega_{\mathrm{as}}$, the adsorbate couples strongly to a surface acoustic mode appearing around $\omega=10\mathrm{cm}^{-1}$. The location of this peak is highly dependent on the dimensions of the simulated surface slab, consistent with the behavior of an acoustic phonon. In Sec. 3, we give a more detailed theoretical and numerical analysis of the size effects observed here, fully accounting for phonon dispersion. In the physisorbed regime, the friction kernel is dominated by this acoustic mode, resulting in highly non-Markovian behavior. The phonon mode associated with this peak is associated with the flexing of the lattice in the direction perpendicular to the surface. Snapshots of this motion are provided in the insets in Fig. 1 and video animations are provided in the electronic SI.
@@
In the chemisorbed regime, the friction kernel is dominated by a high frequency mode around $\omega=210\ \mathrm{cm}^{-1}$. This mode arises from the local oscillations of the surface site the adsorbate is bound to. The frequency of this mode is independent of simulation cell size, suggesting that it is dispersionless. The large amplitude of this mode signifies that in the chemisorbed limit, the adsorbate is primarily sensitive to the local oscillations of the surface binding site (which are shifted in frequency due to the presence of the adsorbate). Many published models for reactive scattering on solid surfaces will describe solid vibrations using only a single, harmonically bound surface atom 6, 39, 40, 41, 42. These results explain why such a method is successful for strongly coupled species.
@@
Despite this large shift in the spectral density when varying $\omega_{\mathrm{as}}$, in Fig 2C we demonstrate that the phonon density of states is nearly identical to that of a bare surface. The only significant change in $\rho(\omega)$ is the presence of the aforementioned surface site local mode present around $\omega=210\ \mathrm{cm}^{-1}$. These results make physical sense given that the low frequency acoustic modes of a solid should be unaffected by gaseous species, especially at low pressures and surface coverages.
@@
To evaluate the effect of the solid force field and structure on these results, we have tested several different crystal structures, facets, elemental compositions, and force fields. The results of these tests are presented in the Section S1 of the SI. While the quantitative properties of the friction kernel vary across different systems (e.g. Ru has a much higher Debye frequency than Pt), the qualitative dependence of the friction kernel on $\omega_{\mathrm{as}}$ is quite general. This independence with respect to the details of the atomistic model is appealing and can be understood from a perturbative perspective. Specifically, in the physisorbed limit, the motion of the surface binding site is a small perturbation to the bulk phonon modes, while in the chemisorbed limit, the bulk phonon modes are a small perturbation to the motion of the surface binding sites. In the SI, we rigorously examine this statement by comparing perturbative schemes to the exact results for $K(t)$ and $\bar{K}(\omega)$. Perturbation theory also allows us to derive analytical approximations to the friction kernel (Eqs. S10-S12) and precisely define the "phase boundaries" in Figure 1.
@@
We acknowledge that in this section we have only studied the effects of phonons for an ideal clean surface. Presumably, sources of surface heterogeneity, such as steps and defects, may add further richness and complexity to the picture we have provided here. We also note that we have only studied elemental solids, and the optical modes of polyatomic crystals could add another interesting dimension to our physical picture. Finally, in this section we have only studied species at a gas-solid interface. Many important surface-chemical processes occur at liquid-solid interfaces, where it has been demonstrated sorption dynamics play a critical role in electrokinetic transport and measurements 43.
@@ 3 Phonon dispersion
In a surface slab, the presence of anisotropy breaks the symmetry of the 3D crystal. This symmetry breaking leads to additional surface mode bands, which can be acoustic (Rayleigh waves), or non-acoustic in character44, 45, 46. In Fig. 4B. and 4C. we demonstrate the phonon dispersion of a Pt(111) surface calculated via EMT. These results were calculated using a 4x4x8 surface replicated in a 6x6 super-cell. We verified convergence with respect to supercell size. Of course, the number of surface phonon bands depends on the size of the surface unit cell used in computing $\mathbf{D}(\mathbf{k})$, however, we demonstrate in in Fig. S7 results are qualitatively similar across different unit cell sizes.
@@
Figure 4: Friction kernel for a model adsorbate on Pt(111) in a finite system (black) and in the infinite system size limit (blue). The finite system is equivalent to the 4x4x8 results presented in Fig. 2.
@@ 4 Phonon effects on reaction rates
One of the primary motivations for studying the atomic vibrations of a solid surface is to elucidate the role of phonons in reactions rates at catalytic interfaces. As was noted first by Kramers47, and expanded on by many others48, 49, 50, 51, the friction of the environment plays a critical role in determining reaction rates, and especially in determining rate prefactors. In this section, we use transition state theory (TST) to demonstrate how surface phonons affect reaction rates, emphasizing the role of the phonon friction kernel. We then specialize to the case of desorption rates and compare the results of our model to experimental temperature dependent rate constants for CO and Xe desorption from Pt(111). Several details of the formalism given this section have been omitted for the purpose of providing a concise presentation of the theory that focuses on physical ramifications. A thorough presentation of the theory is provided in the SI for interested readers.
@@ 4.1.1 General theory
where $N$ is the total number of modes, $f_{i}$ are the vibrational frequencies in the reactant basin, $f_{i}^{{\ddagger}}$ are the non-imaginary vibrational frequencies at the transition state. Note that since the transition state is defined to be a saddle point, it always has one imaginary frequency mode and this mode corresponds the reaction coordinate. In a gas phase molecular reaction, $f_{i}$ is the $i$th eigenfrequency of the molecular mass-weighted Hessian. However, in a reaction at a surface interface, both surface phonon modes and molecular/adsorbate modes will contribute to the rate constant. We thus separate the product of eigenfrequencies into contributions from the solid surface and the molecular adsorbate to yield,
@@
Comparing Eqs. 22 and 23 reveals that phonons introduce two distinct corrections to rate constant: the ratio of solid phonon frequencies between the reactant and transition state, and the shift in adsorbate frequencies. The ratio of phonon frequencies is largest when the reactant molecules are strongly coupled to the surface and the transition state is not. In Fig. 2C we demonstrated that an adsorbate strongly coupled to a surface affects only the highest frequency phonon mode, leaving the bulk of the phonon density of states unchanged. Thus, we can approximate the ratio of phonon frequencies as,
@@
where $\tilde{\omega}_{\mathrm{D}}$ is the highest frequency phonon mode when the reactants are bound to the surface and $\omega_{\mathrm{D}}$ is the bare solid Debye frequency. The factor of $\tilde{\omega}_{\mathrm{D}}$ is directly related to our observation that chemisorbed species couple strongly to local modes above the natural Debye frequency of the metal. The outsized influence of these modes on the dynamics of the chemisorbed species reflects why it appears in the expression for the rate constant. For physisorbed species $\frac{\tilde{\omega}_{\mathrm{D}}}{\omega_{\mathrm{D}}}=1$.
@@
$\Delta\tilde{S}^{\ddagger}=k_{\mathrm{B}}\sum^{N_{A}-1}_{i=10}\ln\left(\frac{\tilde{f}^{{\ddagger}}_{i}}{\tilde{f}_{i}}\right).$ (27)
@@
In Eqs. 23 and 26, the influence of the phonon friction kernel on the rate constant is both explicit and implicit. It is explicit in terms of $K(t=0)$, but also implicit in ratio of normal mode frequencies. The dependence of this ratio on the friction kernel can only be made explicit if additional simplifications are made. Notably, under the condition that the coupling between the adsorbate degrees of freedom and the solid degrees of freedom $\mathbf{C}$ is the same in both the reactant basin and transition state, harmonic TST reduces to Kramers-Grote-Hynes(KGH) theory, where the rate prefactor is explicitly related to the Laplace transform of the friction kernel 52. However, such a condition is not appropriate even in simple desorption processes, and therefore we do not directly use KGH theory here. See the SI (Section S6) for additional details on KGH theory.